# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ D.3.2 Hypothesis test II
D.3.3 Maximum likelihood estimates of the parameters
@@
D.3.4 Details: Statistical details MLE estimate of the variance
@@
D.4 Results: Non-canonical noise scaling is observed in yeast.
To test the RLTO model predictions for noise scaling, we reanalyze the dataset collected by Newman et al., who performed a single-cell proteomic analysis of yeast by measuring the abundance of fluorescent fusions by flow cytometry [24]. Since the competing models (Eqs. 104 and 107) make different scaling predictions, we first apply a statistical test to determine whether the observed scaling is consistent with the canonical model (Eq. 104). We consider the null hypothesis of the canonical model (Eq. 109) and the alternative hypothesis with an unknown scaling exponent (Eq. 110). To test the models, we perform a null hypothesis test. (A detailed description of the statistical analysis, which include the contribution of the noise floor, is given in the Sec. D.1.4.) We reject the null hypothesis with a p-value of $p=6\times 10^{-36}$ The observed scaling exponent is $\hat{a}=-0.57\pm 0.02$, which is close to our predicted estimated exponent from the RLTO model ($-\frac{1}{2}$).
D.5 Methods: Parameter-free prediction of noise from protein-message relation.
@@
D.5.1 Estimating protein number ($\mu_{p}$) for the noise analysis
@@
D.5.2 Empirical models for yeast gene expression
@@
D.5.3 Empirical model for protein number
@@
D.5.4 Empirical model for message number
@@
D.5.5 Empirical model for translation efficiency
@@
Figure S14: Panel A: Three competing models for protein abundance in yeast. The empirical model (purple) fits the slope and the y offset. The RLTO (green) and constant-translation-efficiency (orange) models fit a parameter corresponding to the y offset only. As discussed in the analysis of the proteome fraction, the RLTO model qualitatively captures the scaling of the protein abundance with message number better than the constant translation efficiency model; however, the predicted fit does not correspond to the optimal power law, which is represented by the empirical model. The protein abundance has a cutoff near $10^{1}$ due to autofluorescence [24]. Panel B: Predictions of the noise-protein abundance relation. Using each competing protein abundance model, the noise-protein abundance relation can be predicted using Eq. 106. The canonical noise model (Eq. 104) fails to capture even the scaling of the noise. In contrast, both the RLTO and empirical models quantitatively predict both the scaling and magnitude of the noise. The empirical model has the highest performance, presumably due to its two-parameter fit to the protein abundance in Panel A. A fit accounting for the noise floor is shown in Fig. S12.
D.5.6 Empirical model for the coefficient of variation
@@
D.6 Results: Parameter-free prediction of noise-abundance in yeast
The fit of the competing protein-message abundance models are shown in Fig. LABEL:Mfig:noisefigA. Using each model, we can now predict the relation between protein abundance and noise without additional fitting parameters. The predictions of the three competing models are compared to the experimental data in Fig. LABEL:Mfig:noisefigB.
@@ D.6 Results: Parameter-free prediction of noise-abundance in yeast
D.7 Discussion: Implications of noise
What are the biological implications of gene expression noise? Many important proposals have been made, including bet-hedging strategies, the necessity of feedback in gene regulatory networks, etc. [54]. Our model suggests that robustness to noise fundamentally shapes the central dogma regulatory program. With respect to message number, the one-message-rule sets a lower bound on the transcription rate of essential genes. (See Fig. LABEL:Mfig:principlesB.) With respect to protein expression, robustness to noise has two important implications: Protein overabundance significantly increases protein levels above what would be required in the absence of noise and therefore reshapes the metabolic budget. (See Fig. LABEL:Mfig:principlesA.) Robustness to noise also gives rise to load balancing, the proportionality of the optimal transcription and translation rates. (See Fig. LABEL:Mfig:principlesC.) Not only does robustness to noise affect central dogma regulation, but there is an important reciprocal effect: Load balancing changes the global scaling relation between noise and protein abundance. (See Fig. S14B.)

### v3

@@ D.2.2 Hypothesis test II
D.2.3 Maximum likelihood estimates of the parameters
@@
D.2.4 Details: Statistical details MLE estimate of the variance
@@
D.3 Results: Non-canonical noise scaling is observed in yeast.
To test the RLTO model predictions for noise scaling, we reanalyze the dataset collected by Newman et al., who performed a single-cell proteomic analysis of yeast by measuring the abundance of fluorescent fusions by flow cytometry [8]. Since the competing models (Eqs. 130 and 133) make different scaling predictions, we first apply a statistical test to determine whether the observed scaling is consistent with the canonical model (Eq. 130). We consider the null hypothesis of the canonical model (Eq. 135) and the alternative hypothesis with an unknown scaling exponent (Eq. 136). To test the models, we perform a null hypothesis test. (A detailed description of the statistical analysis, which includes the contribution of the noise floor, is given in the Sec. D.) We reject the null hypothesis with a p-value of $p=6\times 10^{-36}$ The observed scaling exponent is $\hat{a}=-0.57\pm 0.02$, which is close to our predicted estimated exponent from the RLTO model ($-\frac{1}{2}$).
D.4 Methods: Parameter-free prediction of noise from protein-message relation.
@@
D.4.1 Estimating protein number ($\mu_{p}$) for the noise analysis
@@
D.4.2 Empirical models for yeast gene expression
@@
D.4.3 Empirical model for protein number
@@
D.4.4 Empirical model for message number
@@
D.4.5 Empirical model for translation efficiency
@@
Figure S12: Load balancing predicts the scaling of noise. Panel A: Three competing models for protein abundance in yeast. The empirical model (purple) fits the slope and the y offset. The RLTO (green) and constant-translation-efficiency (orange) models fit a parameter corresponding to the y offset only. As discussed in the analysis of the proteome fraction, the RLTO model qualitatively captures the scaling of the protein abundance with message number better than the constant translation efficiency model; however, the predicted fit does not correspond to the optimal power law, which is represented by the empirical model. The protein abundance has a cutoff near $10^{1}$ due to autofluorescence [8]. Panel B: Predictions of the noise-protein abundance relation. Using each competing protein abundance model, the noise-protein abundance relation can be predicted using Eq. 132. The canonical noise model (Eq. 130) fails to capture even the scaling of the noise. In contrast, both the RLTO and empirical models quantitatively predict both the scaling and magnitude of the noise. The empirical model has the highest performance, presumably due to its two-parameter fit to the protein abundance in Panel A. A fit accounting for the noise floor is shown in Fig. S10.
D.4.6 Empirical model for the coefficient of variation
@@
D.5 Results: Parameter-free prediction of noise-abundance in yeast
The fit of the competing protein-message abundance models are shown in Fig. 5A. Using each model, we can now predict the relation between protein abundance and noise without additional fitting parameters. The predictions of the three competing models are compared to the experimental data in Fig. 5B.
@@ D.5 Results: Parameter-free prediction of noise-abundance in yeast
D.6 Discussion: Implications of noise
What are the biological implications of gene expression noise? Many important proposals have been made, including bet-hedging strategies, the necessity of feedback in gene regulatory networks, etc. [7]. Our model suggests that robustness to noise fundamentally shapes the central dogma regulatory program. With respect to message number, the one-message-rule sets a lower bound on the transcription rate of essential genes. (See Fig. 6B.) With respect to protein expression, robustness to noise has two important implications: Protein overabundance significantly increases protein levels above what would be required in the absence of noise and therefore reshapes the metabolic budget. (See Fig. 6A.) Robustness to noise also gives rise to load balancing, the proportionality of the optimal transcription and translation rates. (See Fig. 6C.) Not only does robustness to noise affect central dogma regulation, but there is an important reciprocal effect: Load balancing changes the global scaling relation between noise and protein abundance. (See Fig. S12B.)
Appendix E Data Tables
E.1 Datasets for one-message-rule analysis.
Data S1: Organism: S. cerevisiae (yeast). Original source: [70], [54], [59] & [8]. Essential gene classification: [19]. Processing: We merged these datasets. We added message number (messages per cell-cycle) as described in Sec. C.
Data S2: Organism: Homo sapiens (human). Original sources: mRNA abundances: Human Protein Atlas [60]. Essential gene classification: [51]. Processing: We merged these datasets. We added message number (messages per cell-cycle) as described in Sec. C.
Data S3: Organism: E. coli grown in rich media (LB). Original sources: mRNA abundance: [44]. Essential gene classification: [53]. Processing: We merged these datasets. We added message number (messages per cell-cycle) as described in Sec. C.
Data S4: Organism: E. coli grown in minimal media (MM), Original sources: mRNA abundance: [44]. Essential gene classification: [53]. Processing: We merged these datasets. We added message number (messages per cell-cycle) as described in Sec. C.
E.2 Datasets for load-balancing analysis.
Data S5: Organism: S. cerevisiae (yeast). Original sources: Protein abundance [21]. mRNA abundance: [70]. Processing: We merged these datasets. We added protein number and message number (messages per cell-cycle) as described in Sec. C.
Data S6: Organism: Mus musculus (mammalian). Original source: [23]. Processing: We added protein number and message number (messages per cell-cycle) as described in Sec. C.
Data S7: Organism: E. coli. Original source: [22]. Processing: We added protein number and message number (messages per cell-cycle) as described in Sec. C.
E.3 Datasets for noise analysis.
Data S8: Organism: S. cerevisiae (yeast). Original sources: Noise measurements: [8]. Essential gene classifications: [19]. Protein fluorescence data, rescaled by mass-spec data as described in Sec. D.4.1: [8]. Processing: We merged these datasets. We added protein number as described in Sec. C.